If you start with $1,000 and earn 100% in year 1, lose 50% in year 2, and repeat the year 1 and year 2 returns for 10 years, then what is your average annual return and compound annual return for the 10-year period?
Answer #1
Year Earn
1 1000
2 500
3 1000
4 500
5 1000
6 500
7 1000
8 500
9 1000
10 500

Annual Net Earning = $ 7500

Average Annual Net Earning = $ 7500 / 10 = $ 750

Intitial Investment = $ 1000

Salvage Value = 0

Annual Return = ( 7500 - 1000) *10 /1000 *100 = 65%

Compound Annual Return ( Ending Balance / Inititial Balance ) ^1/n -1

( 7500 /1000) ^1 /10 -1 = 22.32%